OpenStudy (cutiecomittee123):

use the square identities to simplify 2sin^2xsin^2x I tried using sin^2 x =1-cos2x/2 which made the equation into 2(1-cos2x/2)(1-cos2x/2) I am pretty sure I did this wrong. Someone please help and give me some pointers!!!

zepdrix (zepdrix):

\[\large\rm 2\sin^2x \cdot\sin^2x\]This is what they want us to simplify? 0_o

zepdrix (zepdrix):

What do they mean by "simplify"? Do you want only first degree sines and cosines? The directions are very unclear.

OpenStudy (cutiecomittee123):

Well the options for the solutions are 3-4cos(2x)-cos(4x)/4 3+4cos(2x)-cos(4x)/4 3+4cos(2x)+cos(4x)/4 3-4cos(2x)+cos(4x)/4

OpenStudy (cutiecomittee123):

I am just not seeing where any of the numbers are coming from considering we are using the initial equation 2sin^2xsin^2x

zepdrix (zepdrix):

Yah I guess that's what they want then :) Ok your first step looks great! Making use of your Half-Angle Formula:\[\large\rm 2\sin^2x \cdot\sin^2x\quad=2\cdot\frac{1}{2}[1-\cos(2x)]\cdot\frac{1}{2}[1-\cos(2x)]\]

zepdrix (zepdrix):

Let's bring all the fraction business to the front,\[\large\rm =\frac{1}{2}[1-\cos(2x)][1-\cos(2x)]\]We'll have to FOIL this out.

zepdrix (zepdrix):

If we put the cosine stuff in the numerator here it might make things easier for us.\[\large\rm \frac{(1-\cos2x)^2}{2}=\frac{1-2\cos(2x)+\cos^2(2x)}{2}\]Get lost in those steps somewhere? :o What do you lady?

zepdrix (zepdrix):

What do you think? :) typo lol

OpenStudy (cutiecomittee123):

I think I see what has been done. but why include the half angle formula along with the squared identity?

OpenStudy (freckles):

hint use \[\cos^2(u)=\frac{1+\cos(2u)}{2}\] again

OpenStudy (cutiecomittee123):

ohh so we are using the cos2x identity, not sin

OpenStudy (cutiecomittee123):

cos^2x

OpenStudy (freckles):

well you see what @zepdrix has typed... things are in terms cos... you want to write cos^2(2x) in terms of cos(4x) though and that identity I mention will allow you do just that replace u with 2x to see things better

OpenStudy (cutiecomittee123):

okay I am getting it now

zepdrix (zepdrix):

Ya, the first time you had to use the sine^2 identity, then you foil out all the mess, then you apply cosine^2 identity. It's a little tricky ^^

OpenStudy (cutiecomittee123):

okay that is going to be hard to remeber for other equations but practice makes perfect I guess

OpenStudy (cutiecomittee123):

how do I get from here to the solution though?

zepdrix (zepdrix):

Umm lemme write out the steps in a little cleaner way maybe.

OpenStudy (cutiecomittee123):

We have to use cos^2x to solve right? but I dont get how that is plugged in

zepdrix (zepdrix):

\[\large\rm 2\color{orangered}{\sin^2x} \cdot\color{orangered}{\sin^2x}\quad=2\color{orangered}{\left(\frac{1-\cos2x}{2}\right)} \cdot\color{orangered}{\left(\frac{1-\cos2x}{2}\right)}\]Cancel out a pair of 2's, Expand out the numerator,\[\large\rm =\frac{1-2\cos2x+\color{royalblue}{\cos^22x}}{2}\]Applying our Cosine Half-Angle Formula gives us,\[\large\rm =\frac{1-2\cos2x+\color{royalblue}{\left(\dfrac{1+\cos4x}{2}\right)}}{2}\] Oh yes, that step is a little tricky :o I should back up maybe..

zepdrix (zepdrix):

\[\large\rm \cos^2(\color{red}{x})\quad=\frac{1+\cos(2\color{red}{x})}{2}\]What ever angle you start with (in red), you should end up with double that after you apply the formula. So if instead we started with this,\[\large\rm \cos^2(\color{red}{2x})\quad=\frac{1+\cos(2\color{red}{\cdot2x})}{2}\]Do you see why we end up with 4x as the new angle? We're doubling the 2x

OpenStudy (cutiecomittee123):

got it

OpenStudy (cutiecomittee123):

since its called the "double" angle identity

zepdrix (zepdrix):

It's called the Half-Angle Formula :) The angle is cut in half if you read from the other direction. Some teachers reverse the half-angle and double angle formulas... I find that confusing :(

OpenStudy (cutiecomittee123):

okay now what?

OpenStudy (cutiecomittee123):

@zepdrix

zepdrix (zepdrix):

\[\large\rm =\frac{1-2\cos2x+\color{royalblue}{\cos^22x}}{2}\]So you apply your Half-Angle Formula to that cos^2(2x),\[\large\rm =\frac{1-2\cos2x+\color{royalblue}{\left(\dfrac{1+\cos4x}{2}\right)}}{2}\]From here, you `could` get a common denominator and all that fancy business. But I would recommond, instead, that you multiply top and bottom by 2.\[\large\rm =\frac{1-2\cos2x+\left(\dfrac{1+\cos4x}{2}\right)}{2}\cdot\color{orangered}{\frac{2}{2}}\]Which gives us,\[\large\rm =\frac{2-4\cos2x+\left(1+\cos4x\right)}{2}\]And then dro the brackets, nothing is being applied to them. \[\large\rm =\frac{2-4\cos2x+1+\cos4x}{2}\]And combine like-terms. Which is only the 2 and 1 in this case.

zepdrix (zepdrix):

Woops! 4 on the very bottom from multiplying by 2, my bad.\[\large\rm =\frac{2-4\cos2x+1+\cos4x}{4}\]

OpenStudy (cutiecomittee123):

okay and then it would be 3-4cos2x_cos4x/4